# HEIST: A Graph Foundation Model for Spatial Transcriptomics and Proteomics Data

### v1

@@ Abstract
Single-cell transcriptomics has become a great source for data-driven insights into biology, enabling the use of advanced deep learning methods to understand cellular heterogeneity and transcriptional regulation at the single-cell level. With the advent of spatial transcriptomics data we have the promise of learning about cells within a tissue context as it provides both spatial coordinates and transcriptomic readouts. However, existing models either ignore spatial resolution or the gene regulatory information. Gene regulation in cells can change depending on microenvironmental cues from neighboring cells, but existing models neglect gene regulatory patterns with hierarchical dependencies across levels of abstraction. In order to create contextualized representations of cells and genes from spatial transcriptomics data, we introduce HEIST, a hierarchical graph transformer-based foundation model for spatial transcriptomics and proteomics data. HEIST models tissue as spatial cellular neighborhood graphs, and each cell is, in turn, modeled as a gene regulatory network graph. The framework includes a hierarchical graph transformer that performs cross-level message passing and message passing within levels. HEIST is pre-trained on 22.3M cells from 124 tissues across 15 organs using spatially-aware contrastive learning and masked auto-encoding objectives. Unsupervised analysis of HEIST representations of cells, shows that it effectively encodes the microenvironmental influences in cell embeddings, enabling the discovery of spatially-informed subpopulations that prior models fail to differentiate. Further, HEIST achieves state-of-the-art results on four downstream task such as clinical outcome prediction, cell type annotation, gene imputation, and spatially-informed cell clustering across multiple technologies, highlighting the importance of hierarchical modeling and GRN-based representations.
@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) technologies have revolutionized our ability to study gene expression at the resolution of individual cells, offering data-driven insights into biology. The complexity of single-cell data has fueled the development of machine learning [7, 20, 16, 39, 3, 8] methods for modeling cellular diversity, uncovering regulatory mechanisms, predicting cell states, and imputing genes. However, scRNA-seq measurements are unaware of the spatial context of cells within tissues, which is critical for understanding processes such as tissue organization, microenvironment interactions, and how gene regulation gives rise to tissue-level behaviors. Spatial transcriptomics is an emerging technology that bridges this gap by preserving the physical locations of gene expression measurements, enabling comprehensive studies of tissue architecture, cell-cell communication, and tumor microenvironments [42, 31]. Although spatial transcriptomics technology has advanced rapidly, current datasets are small due to low-throughput of sequencing technologies, vary across platforms, organs, and technologies, and often lack detailed labels for supervised learning. These issues requires specialized models to be trained for every new dataset or organ challenging. A foundation model trained on diverse spatial transcriptomics data overcomes these challenges by learning a generalized representation across different tissues, organs, experimental settings, and technologies. This enables application downstream tasks to be performed more effectively, even when data is limited.
While recent foundation models such as scGPT [7] and scFoundation [18] have extended large language models to scRNA-seq data, they do not explicitly model cell-cell relationships. Capturing these cell-cell interactions is vital, as the cellular microenvironment plays a key role in regulating cell states, immune infiltration, tissue development, and tumor progression [22, 6]. In response, graph-based methods such as GraphST [26] and STAGATE [9] have leveraged spatial graphs to model local cell neighborhoods. However, these methods are primarily task-specific and do not support general-purpose, transferable representation learning, limiting their ability to adapt across datasets and organs. CellPLM [39] and scGPT-spatial [7] are foundation models that incorporates cell-cell interactions. But they are pre-trained on a pre-defined set of genes, hindering their generalization to unseen genes. Crucially, none of these approaches capture the influence of gene regulatory networks (GRNs), which drive cell function [28]. GRNs do not act in isolation—they are shaped by, and in turn shape, the spatial microenvironment [49]. These hierarchical relationships between GRNs and cells are fundamental to biological organization [21, 25, 37], ignoring which means missing how gene programs in one cell influence neighboring cells and contribute to tissue-level phenomena [32].
To this end, we propose HEIST (Hierarchical EmbeddIngs for Spatial Transcriptomics), the first spatial transcriptomics foundation model that explicitly models both cell-cell spatial interactions and GRNs, while enabling cross-modal information exchange between cells and genes. In HEIST, cell is associated with a distinct GRN, constructed from co-expression relationships, enabling cross-cell influence between GRNs: as cells interact through spatial graphs, their local microenvironments modulate gene regulation. By explicitly modelling GRNs, HEIST connects gene-level interactions with tissue-level phenotypes, unlocking deeper biological insights that prior models miss. Prior hierarchical graph neural networks [14, 12, 15, 4] either lack mechanisms for cross-level information exchange or only operate on hierarchical graphs within a single data modality (e.g., different levels of resolution in a mesh). HEIST addresses both key challenges of hierarchical modeling: it performs cross-level message passing between genes and cells to integrate spatial and regulatory information. Specifically, gene embeddings are updated using information from their parent cell’s spatial context, while cell embeddings are refined by aggregating their constituent gene representations. As a result, the same gene can acquire different embeddings depending on its spatial context, capturing how microenvironmental factors influence gene regulation, enabling richer representations.
@@
Figure 1: Overview of the HEIST framework. (A) HEIST is pre-trained on a large-scale spatial transcriptomics dataset spanning 124 tissues and 15 organs (22.3M cells). (B) HEIST encodes both gene regulatory networks (GRNs) and spatial cell graphs to support downstream tasks such as cell clustering, gene imputation, and clinical outcome prediction (e.g., immunotherapy response, remission). The HEIST decoder can be fine-tuned while the encoder remains frozen.
@@
We pretrain HEIST (Figure 1(A)) on a large-scale corpus of spatial transcriptomics data comprising over 22.3 million cells from 124 tissues across 15 organs and two technologies. HEIST is evaluated on four downstream tasks (Figure 1(B))—clinical outcome prediction, cell type annotation, gene imputation, and cell clustering—achieving state-of-the-art performance across seven organs. In addition to improved predictive accuracy, HEIST enables the discovery of spatially-informed cellular subpopulations by capturing microenvironmental influences that prior models fail to differentiate. HEIST also offers significant computational efficiency compared to models that learn gene representations, being 8× faster than scGPT-spatial and 48× faster than scFoundation. These results highlight HEIST’s ability to generalize across tissues, technologies, and even unseen genes, without requiring extensive fine-tuning.
We summarize out contributions as follows:
@@
  • Modeling inter-cellular and tissue-scale effects of GRNs: HEIST is the first foundation model for spatial transcriptomics to explicitly incorporate GRNs alongside spatial graphs, enabling a local gene programs to influence tissue-level organization, and vice versa.
@@
  • Hierarchical representation learning with biological inductive bias: Using biologically motivated hierarchical modeling, HEIST captures fine-grained gene regulation within cells and long-range cellular interactions through novel cross-level message passing mechanism, producing adaptive, biologically contextualized embeddings.
@@
  • A task-agnostic, general-purpose foundation model: HEIST is trained in a self-supervised manner on a large-scale corpus of spatial transcriptomics data comprising over 22.3 million cells spanning 15 organs and 124 tissues types. In downstream evaluations, HEIST achieves state-of-the-art performance across four diverse tasks—outperforming prior models, while being computationally efficient.
2 Related works
Hierarchical graph neural networks. Hierarchical graph neural networks operate on graphs with multiple levels, where each level represents a distinct granularity or type of information. MSMGN [12], HOOD [15], and BSMS-GNN [4] perform cross-level message passing across levels, but the levels are of same modality at different resolutions (e.g., coarsened versions of spatial graphs), rather than integrating graphs of distinct modalities and they are geared towards propagating signals globally. HIGH-PPI [14] introduces a hierarchical graph where one level encodes protein-protein interactions and the other encodes residue-level interactions. However, it does not include cross-level message passing, which limits the ability to combine information across biological levels. In contrast, HEIST works information from different modalities while integrating information across levels.
Graph-based Spatial Transcriptomics methods. Graph-based spatial transcriptomics methods construct cell graphs based on spatial proximity and apply graph neural networks to model spatial relationships. SpaGCN [20], CCST [24], STAGATE [9], ConST [48], and GraphST [26] train graph neural networks in self-supervised or unsupervised manner by using techniques such as deep graph infomax (DGI) [36], iterative clustering [2], and auto-encodings task on gene expression [11]. Unlike HEIST, these models focus on specific downstream tasks such as cell clustering and do not generalize to other tissues without retraining. Furthermore, they do not incorporate hierarchical modeling across genes and cells.
scRNA-seq and Spatial Transcriptomics Foundation models. Foundation models for single-cell and spatial transcriptomics aim to pretrain general-purpose representations that can be transferred across datasets and tasks. scFoundation [18], geneformer [33], and scGPT [7] extend transformer architectures to model scRNA-seq data by treating gene expression profiles as sequences. However, these models assume an ordering over genes and do not explicitly capture cell-cell relationships or spatial information. scGPT-spatial [38] adapts scGPT to spatial transcriptomics. CellPLM [39] incorporates cell-cell interactions in pertaining and adds a gaussian mixture prior to overcome data limitations. But CellPLM and scGPT-spatial operate over a fixed set of genes and do not model hierarchy. Existing foundation models thus either neglect spatial dependencies, gene regulatory structure, or both, limiting their ability to capture the context-specific nature of spatial transcriptomics. In contrast, HEIST hierarchically models both this information and computes expressive and generalizable representations by working over GRNs.
@@ 3 Method
In this section, we describe the architecture of HEIST, a hierarchical graph transformer designed to learn multi-level embeddings for spatial transcriptomics data. HEIST takes as input a set of graphs $\{{\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}}),\{{\mathcal{G}}_{g}^{k}({\mathcal{V}},{\mathcal{E}}_{k},{\mathbf{X}}_{k})\}_{k=0}^{|{\mathcal{C}}|}\}$, where ${\mathcal{G}}_{c}$ is a spatial graph capturing cell-cell interactions, ${\mathcal{C}}$ is the set of cells, ${\mathcal{E}}$ is the set of spatial edges, ${\mathbf{P}}\in\mathbb{R}^{|{\mathcal{C}}|\times 2}$ represents the spatial positions, and ${\mathcal{T}}$ is the set of cell types. Each graph ${\mathcal{G}}_{g}^{k}$ represents a GRN within cell $k$, capturing gene-gene interactions specific to that cell; ${\mathcal{V}}$ is the set of genes, and ${\mathcal{E}}_{k}$ and ${\mathbf{X}}_{k}$ denote the edges of GRN and gene expression values for cell $k$, respectively. We explain the graph-creation process in Appendix Section A. Given these graphs, HEIST computes cell and gene embeddings ${\mathbf{Z}}_{c}\in\mathbb{R}^{|{\mathcal{C}}|\times d}$ and ${\mathbf{Z}}_{g}\in\mathbb{R}^{|{\mathcal{C}}||{\mathcal{V}}|\times d}$, such that
${\mathbf{Z}}_{c},{\mathbf{Z}}_{g}=\text{HEIST}\left({\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}}),\{{\mathcal{G}}_{g}^{k}({\mathcal{V}},{\mathcal{E}}_{k},{\mathbf{X}}_{k})\}_{k=0}^{|{\mathcal{C}}|}\right).$
The model first computes positional embeddings. As shown in Figure 2, model performs intra-level message passing within each graph, followed by cross-level message passing to integrate multi-modal information. HEIST is pre-trained using a combination of contrastive and auto-encoding objectives on gene expression and cell locations. By using these components, HEIST can learn expressive and context-aware cell and gene embeddings that reflect biologically meaningful relationships between cells and genes.
@@
The input cell embeddings ${\mathbf{H}}_{c}^{(0)}$ and gene embeddings for cell $k$ ${\mathbf{H}}_{g}^{k^{(0)}}$ are passed through HEISTLayer $L$ times, and the representations are calculated using the equation below:
${\mathbf{H}}_{c}^{(l)},{\mathbf{H}}_{g}^{(l)}=\texttt{HEISTLayer}({\mathbf{H}}_{c}^{(l-1)},{\mathbf{H}}_{g}^{(l-1)},{\mathcal{E}},\{{\mathcal{E}}_{k}\}_{k=0}^{|{\mathcal{C}}|})$
@@
$\tilde{{\mathbf{H}}}_{c}^{(l)}=\texttt{CellGraphTransformer}({\mathbf{H}}_{c}^{(l-1)},{\mathcal{E}}),\tilde{{\mathbf{H}}}_{g}^{(l)}=\texttt{GeneGraphTransformer}({\mathbf{H}}_{g}^{(l-1)},\{{\mathcal{E}}_{k}\}_{k=0}^{|{\mathcal{C}}|})$
where CellGraphTransformer and GeneGraphTransformer are graph transformers as explained in the Appendix Section B.2.
Cross-level message passing. To integrate spatial and gene modalities, HEIST performs cross-level message passing between cell and gene graphs at each layer (Figure 2). Gene embeddings are updated based on spatial context through their parent cell’s embedding, while cell embeddings are refined using pooled summaries of their genes, ensuring transcriptional states shape spatial identity. This bidirectional interaction captures tissue hierarchy, where gene expression depends on both regulatory signals and spatial microenvironments. HEIST thus learns representations that reflect both local gene regulation and large-scale tissue structures.
We use a directional attention mechanism to perform cross-level message passing between gene and cell graphs. Let $\tilde{{\mathbf{H}}}_{g}^{(l)}$ and $\tilde{{\mathbf{H}}}_{c}^{(l)}$ denote the intermediate gene and cell embeddings at layer $l$ after intra-level updates and before cross-level integration. HEIST updates these representations using the equation below:
@@
where $\tilde{{\mathbf{H}}}_{c}^{(l)^{\text{repeat}}}\in\mathbb{R}^{|{\mathcal{C}}||{\mathcal{V}}|\times d}$ represents the cell embeddings repeated $|{\mathcal{V}}|$ times to align with the gene embeddings in each cell. Each gene receives information from its parent cell, enabling spatial context to modulate gene-level representations. Conversely, $\bar{{\mathbf{H}}}_{g}^{(l)}\in\mathbb{R}^{|{\mathcal{C}}|\times d}$ is obtained by aggregating the gene embeddings within each cell: $\bar{{\mathbf{H}}}_{g}^{(l)}=[\texttt{AGG}(\tilde{{\mathbf{H}}}_{g}^{1}),\ldots,\texttt{AGG}(\tilde{{\mathbf{H}}}_{g}^{|{\mathcal{C}}|})]$, where $\texttt{AGG}(\cdot)$ can be aggregation function such as MEAN pooling or differential pooling [45]. This allows the cell embedding to be informed by the internal transcriptional state of the cell. We use the directional attention mechanism to perform cross message passing as shown in equation below:
[…]
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ are learned weight matrices, and <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > is the inner product computed row-wise.
Advantages. This formulation enables targeted, direction-aware communication between modalities while preserving their structure and semantics. It is well suited for spatial transcriptomics, where cell and gene representations must be coupled yet retain distinct meanings. Gene expression depends not only on intrinsic regulation but also on spatial context, such as cell interactions and microenvironments. By letting each gene attend to its parent cell embedding, HEIST shapes gene representations in a cell-specific, spatially informed way, and vice versa. Unlike symmetric attention or feature concatenation, this directional mechanism preserves data hierarchy, respects modality roles, and captures how local gene regulation drives tissue organization. Because directional message passing respects the natural hierarchy between genes and cells, allowing each to influence the other without collapsing their distinct biological roles, enabling HEIST to learn spatially informed, biologically grounded representations that generalize across tissues and experiments.
@@
${\mathbf{Z}}_{c},{\mathbf{Z}}_{g}=\text{HEIST}\left({\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}}),\{{\mathcal{G}}_{g}^{k}({\mathcal{V}},{\mathcal{E}}_{k},{\mathbf{X}}_{k})\}_{k=0}^{|{\mathcal{C}}|}\right).$
@@
$\hat{{\mathbf{P}}},\{\hat{{\mathbf{X}}}_{k}\}_{k=0}^{|{\mathcal{C}}|}=\text{HEIST-Decoder}\left({\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}}),\{{\mathcal{G}}_{g}^{k}({\mathcal{V}},{\mathcal{E}}_{k})\}_{k=0}^{|{\mathcal{C}}|},{\mathbf{Z}}_{c},{\mathbf{Z}}_{g}\right).$
where, HEIST-Decoder is a 3-layer GIN network [43].

### v2

@@ Abstract
Single-cell transcriptomics and proteomics have become a great source for data-driven insights into biology, enabling the use of advanced deep learning methods to understand cellular heterogeneity and gene expression at the single-cell level. With the advent of spatial-omics data, we have the promise of characterizing cells within their tissue context as it provides both spatial coordinates and intra-cellular transcriptional or protein counts. Beyond transcriptomics, proteomics offers a complementary view by directly measuring proteins, which are the primary effectors of cellular function and key therapeutic targets. However, existing models either ignore the spatial information or the complex genetic and proteomic programs within cells. Thus they cannot infer how cell internal regulation adapts to microenvironmental cues. Furthermore, these models often utilize fixed gene vocabularies, hindering their generalizability to datasets with different genes than pretraining. In this paper, we introduce HEIST, a hierarchical graph transformer foundation model for spatial transcriptomics and proteomics. HEIST models tissues as hierarchical graphs. The higher level graph is a spatial cell graph, and each cell in turn, is represented by its lower level gene co-expression network graph. Rather than using a fixed gene vocabulary, HEIST computes gene embeddings from its co-expression network and cellular context. HEIST achieves this by performing both intra-level and cross-level message passing to utilize the hierarchy in its embeddings and can thus generalize to novel datatypes including spatial proteomics without retraining. HEIST is pretrained on 22.3M cells from 124 tissues across 15 organs using spatially-aware contrastive and masked autoencoding objectives. Unsupervised analysis of HEIST embeddings reveals spatially informed subpopulations missed by prior models. Downstream evaluations demonstrate generalizability to proteomics data and state-of-the-art performance in clinical outcome prediction, cell type annotation, and gene imputation across multiple technologies.
@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) has revolutionized our ability to study gene expression at the resolution of individual cells, offering data-driven insights into biology. The complexity of these datasets has fueled the development of machine learning methods for modeling cellular diversity, predicting cell states, and imputing or denoising gene expression values  (Cui et al., 2024; Hu et al., 2021; Grønbech et al., 2020; Wen et al., 2023a; Bravo Gonzalez-Blas et al., 2024; Dijk et al., 2017). However, a limitation of scRNA-seq is the lack of spatial context of cells within tissues, which is important for understanding processes such as tissue organization, microenvironment interactions and how gene co-expression influences tissue-level behaviors. Spatial transcriptomics is an emerging technology that bridges this gap by preserving the physical locations of gene expression measurements, enabling a holistic study of tissue architecture, cell-cell communication, and their aberrations in the tumor context (Xiao & Yu, 2021; Rodriques et al., 2019). Similarly, spatial proteomics assays were developed to directly capture protein abundance, expression and signaling pathways, offering a complementary layer of biological insight. Despite advances in spatial omics, datasets remain limited by low throughput, platform and tissue heterogeneity, and scarce labels, often necessitating dataset-specific models (e.g., MIBI (Angelo et al., 2014), Imaging CyTOF). A foundation model trained on diverse spatial omics can address these challenges by learning generalized representations across tissues, organs, settings, and technologies, enabling strong performance on downstream tasks even with limited data.
To this end, we propose HEIST (Hierarchical EmbeddIngs for Spatial Transcriptomics), the first foundation model for spatial transcriptomics that explicitly models both spatial proximity and internal co-expression networks while also enabling cross-modal transfer to proteomics. Through cross-level message passing, internal embeddings are shaped by the spatial context of their parent cell, while cell embeddings are updated from their constituent genes. This produces adaptive representations, allowing the same gene to be encoded differently depending on context. In this way, HEIST unifies spatial and molecular hierarchies, linking gene-level interactions to tissue-level phenotypes.
Prior foundation models such as scGPT (Cui et al., 2024), scFoundation (Hao et al., 2023), and CellPLM (Wen et al., 2023a) either neglect cell–cell structure or are limited to predefined gene sets, hindering generalization to unseen genes and proteins. Graph-based methods like GraphST (Long et al., 2023) and STAGATE (Dong & Zhang, 2022) (Explained in detail in Appendix A) capture spatial neighborhoods but remain task-specific and non-transferable. While transformer-based models like scGPT-spatial (Wang et al., 2025) treat all genes as fully connected graph and fail to model the inductive bias of co-expression. HEIST overcomes these gaps by associating each cell with a co-expression network that interacts with its spatial neighbors, incorporating microenvironmental cues into representations.
@@
Figure 1: Overview of the HEIST framework. (A) HEIST is pre-trained on a large-scale spatial transcriptomics dataset spanning 124 tissues and 15 organs (22.3M cells). (B) HEIST encodes both gene co-expression networks and spatial cell graphs to support downstream tasks such as cell clustering, gene imputation, and clinical outcome prediction (e.g., immunotherapy response, remission). The HEIST decoder can be fine-tuned while the encoder remains frozen.
@@
We pretrain HEIST (Figure 1(A)) on 22.3 million cells from 124 tissues across 15 organs and two technologies, and evaluate on four downstream tasks (Figure 1(B)): clinical outcome prediction, cell type annotation, gene imputation, and cell clustering. We achieve state-of-the-art performance across seven organs. HEIST enables the discovery of spatially-informed cellular subpopulations that previous models fail to do and is 8× faster than scGPT-spatial and 48× faster than scFoundation. We summarize our contributions as follows:
@@
  • Modeling inter-cellular and hierarchical effects of co-expression networks: HEIST is the first foundation model for spatial omics to explicitly incorporate co-expression networks alongside spatial graphs in a hierarchical graph, enabling a local gene programs to influence tissue-level organization, and vice versa.
@@
  • Hierarchical representation learning with biological inductive bias: Using biologically motivated hierarchical modeling, HEIST captures fine-grained gene co-expression within cells and long-range cellular interactions through novel cross-level message passing, producing biologically contextualized embeddings.
@@
  • A task-agnostic, general-purpose foundation model: HEIST is trained in a self-supervised manner on a large-scale corpus of spatial transcriptomics data comprising over 22.3M cells spanning 15 organs and 124 tissues. In downstream evaluations, HEIST achieves state-of-the-art performance across four diverse tasks—outperforming prior models, generalizes to proteomics data, while being computationally efficient.
@@ 2 Method
In this section, we describe the architecture of HEIST, a hierarchical graph transformer designed to learn multi-level embeddings for spatial transcriptomics and proteomics data. HEIST takes as input, a set of graphs $\{{\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}}),\{{\mathcal{G}}_{g}^{t_{k}}({\mathcal{V}},{\mathcal{E}}_{t_{k}},{\mathbf{X}}_{k})\}_{k=0}^{|{\mathcal{C}}|}\}$, where ${\mathcal{G}}_{c}$ is a spatial graph capturing spatial proximity, $\mathcal{C}$} is the set of cells, $\mathcal{E}$} is the set of spatial edges, ${\mathbf{P}}\in\mathbb{R}^{|{\mathcal{C}}|\times 2}$ represents the spatial positions, $\mathcal{T}$} is the set of cell types, and $t_{k}$} is cell-type of cell $k$. Each graph ${\mathcal{G}}_{g}^{t_{k}}$ represents a gene co-expression network within cell $k$, $\mathcal{V}$} is the set of genes, and ${\mathcal{E}}_{t_{k}}$ and ${\mathbf{X}}_{k}$ denote the edges of the gene co-expression network and expression values for cell $k$, respectively.
Hierarchical Graph Construction. As shown in Figure A1, we first preprocess the data by removing outliers, normalizing gene expression, and retaining highly variable genes. Then we apply MAGIC (Dijk et al., 2017) to denoise gene expression values and reduce dropout noise. To build the gene co-expression networks, we first subset the cells based on cell-types using provided annotations or Leiden clustering. Following this, we compute pairwise mutual information between denoised genes within each type, and connect gene pairs above a threshold $\tau$. This results in total of $|{\mathcal{T}}|$ gene co-expression networks with mutual information prior. We create the spatial cell–cell graph by computing Voronoi polygons from cell coordinates and connecting cells in adjacent polygons. We then connect each cell with the gene co-expression network of that cell-type. The resulting outputs are a spatial graph ${\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}})$ and a set of gene co-expression networks ${{\mathcal{G}}_{g}^{t_{k}}({\mathcal{V}},{\mathcal{E}}_{t_{k}},{\mathbf{X}}_{k})}_{k=1}^{|{\mathcal{C}}|}$. We discuss the graph creation at length in Appendix B. Given these graphs, HEIST computes cell embeddings and gene embeddings ${\mathbf{Z}}_{c}\in\mathbb{R}^{|{\mathcal{C}}|\times d}$ and ${\mathbf{Z}}_{g}\in\mathbb{R}^{|{\mathcal{C}}||{\mathcal{V}}|\times d}$, such that
${\mathbf{Z}}_{c},{\mathbf{Z}}_{g}=\text{HEIST}\left({\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}}),\{{\mathcal{G}}_{g}^{t_{k}}({\mathcal{V}},{\mathcal{E}}_{t_{k}},{\mathbf{X}}_{k})\}_{k=0}^{|{\mathcal{C}}|}\right).$
As shown in Figure 2, the model first performs intra-level message passing (Equation 1) within each graph, followed by cross-level message passing (Equation 2) to integrate multi-modal information. HEIST is pre-trained using a combination of contrastive and auto-encoding objectives on gene expression and cell locations. By using these components, HEIST can learn expressive and context-aware cell and gene embeddings that reflect biologically meaningful relationships between cells and genes. Note that as a result of this setup, gene representations are themselves learned in the context of the hierarchical graph, instead of based on a fixed gene vocabulary. They are initialized with rank-based and sinusoidal positional encodings, and dynamically updated through message passing in co-expression graphs, allowing HEIST to generalize to unseen genes or proteomic features by grounding embeddings in co-expression dynamics.
@@
First we initialize the input cell embeddings ${\mathbf{H}}_{c}^{(0)}$ and gene embeddings for cell $k$ ${\mathbf{H}}_{g}^{k^{(0)}}$ using positional encodings explained in Appendix C.2. Then they are passed through HEISTLayer $L$ times, and the representations are calculated using the equation below:
${\mathbf{H}}_{c}^{(l)},{\mathbf{H}}_{g}^{(l)}=\texttt{HEISTLayer}({\mathbf{H}}_{c}^{(l-1)},{\mathbf{H}}_{g}^{(l-1)},{\mathcal{E}},\{{\mathcal{E}}_{t_{k}}\}_{k=0}^{|{\mathcal{C}}|})$
@@
$\tilde{{\mathbf{H}}}_{c}^{(l)}=\texttt{CellGraphTransformer}({\mathbf{H}}_{c}^{(l-1)},{\mathcal{E}}),\tilde{{\mathbf{H}}}_{g}^{(l)}=\texttt{GeneGraphTransformer}({\mathbf{H}}_{g}^{(l-1)},\{{\mathcal{E}}_{t_{k}}\}_{k=0}^{|{\mathcal{C}}|})$ (1)
where CellGraphTransformer and GeneGraphTransformer are graph transformers as explained in the Appendix Section C.2.
Cross-level message passing. To integrate spatial and gene modalities, HEIST performs cross-level message passing between cell and gene graphs at each layer (Figure 2). Gene embeddings are updated based on spatial context through their parent cell’s embedding, while cell embeddings are refined using pooled summaries of their genes, ensuring transcriptional states shape spatial identity. This bidirectional interaction captures tissue hierarchy, where gene expression depends on both co-expression signals and spatial microenvironments. HEIST thus learns representations that reflect both local gene co-expression and large-scale tissue structures.
We use the directional attention mechanism to perform cross message passing as shown in equation below:
$\texttt{CrossMessagePassingLayer}({\mathbf{H}}_{\text{to}},{\mathbf{H}}_{\text{from}})=\left(\frac{<{\mathbf{H}}_{\text{to}}\mathbf{W}_{q},{\mathbf{H}}_{\text{from}}\mathbf{W}_{k}>}{\sqrt{d}}\right)\cdot({\mathbf{H}}_{\text{to}}\mathbf{W}_{v}),$ (2)
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$ are learned weight matrices, and $\cdot$$\cdot$> is the row-wise inner product. Let $\tilde{{\mathbf{H}}}_{g}^{(l)}$ and $\tilde{{\mathbf{H}}}_{c}^{(l)}$ denote the intermediate gene and cell embeddings at layer $l$ after intra-level updates and before cross-level integration. HEIST updates these representations using the equation below:
@@
where $\tilde{{\mathbf{H}}}_{c}^{(l)^{\text{repeat}}}\in\mathbb{R}^{|{\mathcal{C}}||{\mathcal{V}}|\times d}$ represents the cell embeddings repeated $|{\mathcal{V}}|$ times to align with the gene embeddings in each cell. Each gene receives information from its parent cell, enabling spatial context to modulate gene-level representations. Conversely, $\bar{{\mathbf{H}}}_{g}^{(l)}\in\mathbb{R}^{|{\mathcal{C}}|\times d}$ is obtained by aggregating the gene embeddings within each cell: $\bar{{\mathbf{H}}}_{g}^{(l)}=[\texttt{AGG}(\tilde{{\mathbf{H}}}_{g}^{1}),\ldots,\texttt{AGG}(\tilde{{\mathbf{H}}}_{g}^{|{\mathcal{C}}|})]$, where $\texttt{AGG}(\cdot)$ can be aggregation function such as MEAN pooling or differential pooling (Ying et al., 2018). This allows the cell embedding to be informed by the internal transcriptional state of the cell.
Advantages. This formulation enables targeted, direction-aware communication between modalities while preserving their structure and semantics. It is well suited for spatial transcriptomics and proteomics data, where cell and gene representations must be coupled yet retain distinct meanings. By letting each gene attend to its parent cell embedding, HEIST shapes gene representations in a cell-specific, spatially informed way, and vice versa. Unlike symmetric attention or feature concatenation, this directional mechanism preserves data hierarchy, respects modality roles, and captures how local gene co-expression drives tissue organization. Directional message passing respects the natural hierarchy between genes and cells, allowing each to influence the other without collapsing their distinct biological roles, enabling HEIST to learn spatially informed, biologically grounded representations that generalize across tissues and proteomics.
@@
${\mathbf{Z}}_{c},{\mathbf{Z}}_{g}=\text{HEIST}\left({\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}},{\mathbf{P}},{\mathcal{T}}),\{{\mathcal{G}}_{g}^{t_{k}}({\mathcal{V}},{\mathcal{E}}_{t_{k}},{\mathbf{X}}_{k})\}_{k=0}^{|{\mathcal{C}}|}\right).$
@@
$\hat{{\mathbf{P}}},\{\hat{{\mathbf{X}}}_{k}\}_{k=0}^{|{\mathcal{C}}|}=\text{HEIST-Decoder}\left({\mathcal{G}}_{c}({\mathcal{C}},{\mathcal{E}}),\{{\mathcal{G}}_{g}^{t_{k}}({\mathcal{V}},{\mathcal{E}}_{t_{k}})\}_{k=0}^{|{\mathcal{C}}|},{\mathbf{Z}}_{c},{\mathbf{Z}}_{g}\right).$
where, HEIST-Decoder is a 3-layer GIN network (Xu et al., 2018).